40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5261     Entry time: Thu Aug 18 10:17:04 2011
Author: kiwamu, steve 
Type: Update 
Category: SUS 
Subject: oplevs reestablished at Vertex 

Kiwamu and Steve, from yesterday

PRM and BS oplev paths were relaid after setting 1/2 OSEM voltages. The incident beam on suspended optics are centered  within  ~ +- 2 mm

I noticed many unvected ss screws are used on the big Al table tops. The SS 1/4-20 screws

 used on the optical tables in vacuum MUST be VENTED!

Also, please use  SS clamps. Replace aluminum ones when you can. We have plenty baked ones.

 

 

ELOG V3.1.3-